OpenStudy (nikolas):

Prove that 1/(1+cosx) = csc^2(x) - cotx/sinx I've tried a few things and can't seem to get this right... I think the first step is to change csc^2x into (1+cot^2x) and then get a common denominator of sinx, but either this is the wrong way or I'm making more mistakes further on... any ideas? Thanks.

OpenStudy (anonymous):

well your idea is absolutely right actually \[ \large \frac{1}{\sin^2(x)}- \frac{\cot x \sin x}{\sin ^2x}=\frac{1-\cos(x)}{1-\cos^2x } \\ \frac{1-\cos(x)}{1-\cos^2x} =\frac{1-\cos x}{(1+\cos x)(1-\cos x) }= \frac{1}{1+\cos(x)}\]

hero (hero):

0. Given \(\large\frac{1}{1+\cos x} = \csc ^2x - \frac{\cot x}{\sin x}\) 1. Replace csc^2x with 1/sin^2x: \(\large\frac{1}{1+\cos x} = \frac{1}{\sin^2x} - \frac{\cot x}{\sin x}\) 2. Rewrite cotx/sinx as 1/sinx times cos x/sin x: \(\large\frac{1}{1+\cos x} = \frac{1}{\sin^2x} - \frac{1}{\sin x} \dot\ \frac{\cos x}{\sin x}\) 3. Multiply: \(\large\frac{1}{1+\cos x} = \frac{1}{\sin^2x} - \frac{\cos x}{\sin^2x}\) 4. Combine fractions: \(\large\frac{1}{1+\cos x} = \frac{1-cos x}{\sin^2x}\) 5. Cross Multiply: \(\sin^2x = (1+\cos x)(1-\cos x)\) 6. Muliply the right side to get: \(\sin^2x = 1 - \cos^2 x\) 7. Add cos^2x to both sides to get: \(\sin^2x + cos^2x = 1\) 8. Simplify: \(1 = 1\) 9. Solution: True

OpenStudy (nikolas):

Thank you both! I ended up getting a common denominator of sin^2x(1+cosx), converting the numerator to sin^2x and cancelling it out to get 1/(1+cosx).